OpenStudy (quickstudent):

Which equation has the solution set x = {2,3} ? a. 2/3x = 0 b. (x - 2) (x - 3) = 0 c. (x + 2) (x + 3) = 0 d. (2x + 3)^2 = 0

OpenStudy (quickstudent):

I think it's b.

OpenStudy (radar):

Look at your choices. a. Is an impossibility unless 2 = 0 c. would result in x= -2 and x = -3 d would result in a double root x=-3

OpenStudy (radar):

I do too.

OpenStudy (quickstudent):

Thanks!

jimthompson5910 (jim_thompson5910):

It's always a good idea to check any possible solutions ------------------------------- Checking x = 2 (x - 2)(x - 3) = 0 (2 - 2)(2 - 3) = 0 ... plug in x = 2 (0)(-1) = 0 0 = 0 ... works ------------------------------- Checking x = 3 (x - 2)(x - 3) = 0 (3 - 2)(3 - 3) = 0 ... plug in x = 3 (1)(0) = 0 0 = 0 ... works ------------------------------- Both possible solutions x = 2 and x = 3 work